# MetaCluster: Enabling Deep Compression of Kolmogorov-Arnold Network

### v1

@@ 4.2 Results
4.2.1 Fully-Connected Network
Table 1 demonstrates the high classification accuracy and compression factor of fully-connected MetaCluster models across all datasets and basis functions. For example, in the case of MetaClusterKAN, we can see that on Cifar-10, it achieves a compression factor of up to $10.7\times$ (from 410.82KB to 38.32KB) compared to the MetaKAN, and up to $79.9\times$ (from 3064.95 KB to 38.32KB) compared to the KAN. Furthermore, it achieves this compression factor while matching the classification accuracy of MetaKAN. Although the Meta variants are capable of achieving high accuracy without fine-tuning, this is not the case for the non-Meta variants. For example, in the case of GramKAN, it experiences a decrease in classification accuracy of up to $3.9\times$ (from 21.92% to 5.61%), which, when recovered with fine-tuning, remains a $1.5\times$ decrease in accuracy (from 21.92% to 14.51 %). Such results showcase the importance of leveraging a meta-learner to learn a set of activations that sit on a lower-dimensional subspace to aid in classification accuracy after clustering in fully-connected models.
@@ 4.2.1 Fully-Connected Network
KAN 1,031.44 95.36 3,064.95 47.52 3,177.45 17.90
MetaKAN 141.32 96.45 410.82 46.99 422.08 20.08
ClusterKAN 13.84 63.47 38.34 29.82 39.75 7.32
   + fine-tune 13.84 93.34 38.34 43.73 39.75 12.54
MetaClusterKAN 13.84 96.02 38.34 46.39 39.75 19.03
   + fine-tune 13.84 96.45 38.34 46.59 39.75 19.43
FastKAN 900.56 96.75 2,676.83 49.28 2,778.08 20.76
MetaFastKAN 108.09 96.29 316.35 47.41 327.60 19.89
ClusterFastKAN 20.54 32.85 57.30 23.53 58.71 7.63
   + fine-tune 20.54 87.37 57.30 38.26 58.71 12.85
MetaClusterFastKAN 20.54 96.28 57.30 47.46 58.71 19.65
   + fine-tune 20.54 96.58 57.30 47.35 58.71 19.54
GramKAN 497.46 96.11 1,477.48 51.15 1,534.43 21.92
MetaGramKAN 101.49 95.83 297.49 49.31 309.45 19.74
ClusterGramKAN 13.96 48.16 38.46 30.32 40.58 5.61
   + fine-tune 13.96 85.22 38.46 45.36 40.58 14.51
MetaClusterGramKAN 14.15 95.22 38.65 49.06 40.76 18.87
   + fine-tune 14.15 95.82 38.65 49.31 40.76 19.46
4.2.2 Convolutional Network
Table 2 demonstrates that, aside from MetaFastKANConv, the Meta variants match or exceed the classification accuracy of all the non-Meta variants at a reduced memory cost. For example, in the case of MetaKANConv on Cifar-10, the reduced memory cost is up to $4.5\times$ (from 13.7 MB to 3.1 MB) compared to KANConv. Upon clustering, we find that the compression factor increases by a factor of $7.1\times$ (from 3.1 MB to 435.34 KB). In total, compared to the KANConv, the MetaClusterKANConv achieves a reduction of up to $31.7\times$ (from 13.7 MB to 435.34 KB). For such a reduction in storage cost, after fine-tuning, we find there is a negligible impact on the downstream classification accuracy when comparing KANConv to MetaKANConv.
Although the non-Meta variants recover accuracy from the fine-tuning step, the recovery still leaves them far behind their original state. For instance, upon clustering, FastKANConv experiences up to a $5.3\times$ decrease (from 76.06% to 14.24 % ) in classification accuracy. Then, when recovered with fine-tuning, it still retains a $1.4\times$ decrease (from 76.06% to 55.14 %) in classification accuracy. Once again, as with the fully-connected architecture, these results demonstrate the importance of our MetaCluster framework for maintaining accuracy with a high compression factor.
@@ 4.2.2 Convolutional Network
KANConv 13,634.00 99.55 13,654.27 73.15 13,744.62 21.46
MetaKANConv 3,048.40 99.25 3,052.90 72.45 3,143.25 43.59
ClusterKANConv 429.95 94.81 430.51 33.98 520.87 11.32
   + fine-tune 429.95 99.35 430.51 57.59 520.87 20.85
MetaClusterKANConv 434.77 99.21 435.34 67.50 525.70 37.06
   + fine-tune 434.77 99.11 435.34 71.67 525.70 43.31
FastKANConv 13,632.09 99.27 13,652.37 76.06 13,742.72 45.40
MetaFastKANConv 3,041.57 99.06 3,046.08 71.42 3,136.43 41.97
ClusterFastKANConv 428.03 72.87 428.61 14.24 518.97 6.47
   + fine-tune 428.03 97.35 428.61 55.14 518.97 23.59
MetaClusterFastKANConv 427.97 99.07 428.55 71.06 518.91 41.39
   + fine-tune 427.97 99.06 428.55 71.38 518.91 41.60
GramKANConv 7,586.47 99.25 7,597.72 80.92 7,688.07 47.10
MetaGramKANConv 3,047.93 99.50 3,052.43 81.88 3,142.79 53.26
ClusterGramKANConv 418.93 84.83 419.49 12.44 509.85 2.39
   + fine-tune 418.93 99.37 419.49 76.35 509.85 43.77
MetaClusterGramKANConv 418.81 99.48 419.38 79.36 509.73 51.56
   + fine-tune 418.81 99.53 419.38 81.37 509.73 52.63
@@ 4.3.1 Basis Coefficient Count
We investigate the impact of the basis coefficient count (i.e., the number of radial basis functions) on downstream clustering performance. Since our technique employs a meta-learner to learn a lower-dimensional subspace, our approach should be resilient to changes in the basis coefficients we cluster. From Table 3, we can see that the initial accuracy of MetaFastKAN and MetaFastKANConv accuracy peaks at a coefficient count of 5 before dropping. However, even with these variations in accuracy for the coefficient count, the relative percentage change remains the same. This indicates that, regardless of the coefficient count, our MetaCluster framework remains capable of learning a lower-dimensional manifold, which facilitates easier clustering of weights.
Table 3: Impact of basis coefficient count on memory footprint and classification accuracy.
Coeff count Memory Acc% Memory Acc% %Chg Memory Acc% Memory Acc% %Chg
5 316.35 47.41 57.30 47.46 0.11 3,046.08 71.42 428.55 71.06 -0.50
8 316.76 47.62 57.68 46.96 -1.39 3,046.52 69.61 440.61 68.66 -1.36
10 317.03 46.91 57.93 47.04 0.28 3,046.81 66.78 448.64 66.58 -0.30
15 317.71 45.88 58.55 45.15 -1.59 3,047.54 64.24 468.72 63.99 -0.39
20 318.40 43.57 59.18 43.75 0.41 3,048.26 54.41 488.80 54.45 0.07
@@ 4.3.2 Cluster Count
We verify that the MetaCluster framework scales well with the cluster count by comparing its cluster scaling properties with those of a non-meta KAN. As shown in Figure 2(b), MetaCluster maintains accuracy well as clusters are reduced, especially after fine-tuning. For the convolutional model, decreasing the cluster count from 256 to 32 lowers MetaClusterFastKANConv accuracy by only $2\%$, whereas ClusterFastKANConv drops by $21.39\%$ over the same range. Such results demonstrate that our approach has the capability of achieving even greater levels of compression, albeit at the expense of a minor accuracy impact.
@@ 4.3.3 Meta-learner Embedding Size
The embedding size of the metalearner directly influences the ease of clustering the meta-learner generated weights. The general trend is that as the embedding dimension increases, it becomes increasingly difficult to cluster the generated weights appropriately. While we have qualitatively demonstrated this in Figure 1, where we see that increasing embedding dimension causes the plot to become less structured, we provide a quantitative analysis of the effect in Appendix A.2.2.
@@ 5.1 Kolmogorov-Arnold Networks
The KAN has emerged as a popular alternative to the MLP (Liu et al., 2024). The most popular variant has used a weighted summation of a B-spline basis to represent the activations on each edge (Liu et al., 2024). However, since its rise to popularity, there have been numerous alternative variations of the KAN, such as using a host of basis functions, such as RBFs, Chebyshev polynomials, Legendre Polynomials, Gram polynomials, wavelets, and rational functions (Li, 2024; SS et al., 2024; Bozorgasl & Chen, 2024; Aghaei, 2024).
Although KAN has primarily achieved success in scientific applications and equation modeling tasks (Li et al., 2025; Wang et al., 2024; 2025; Coffman & Chen, 2025; Koenig et al., 2024), it has recently shown promise in the realm of computer vision with the introduction of the Kolmogorov-Arnold Transformer (Yang & Wang, 2024; Raffel & Chen, 2025). However, despite its successes, the KAN has struggled with shortcomings such as increased training instability, computational cost, and parameter count (Yu et al., 2024; Chen et al., 2024). Our work directly targets the memory scaling obstacle of KANs, creating topologically identical KANs that are smaller than comparable MLPs while maintaining accuracy.

### v2

@@ 4.1 Image Classification Results
4.1.1 Fully-Connected Image Classification
@@
KAN 1,031.44 $95.37\pm 0.11$ 3,064.95 $47.52\pm 0.16$ 3,177.45 $17.80\pm 0.10$
MetaKAN 141.32 $96.36\pm 0.09$ 410.82 $46.28\pm 0.39$ 422.08 $20.17\pm 0.24$
ClusterKAN 13.84 $61.37\pm 2.77$ 38.34 $27.92\pm 2.88$ 39.75 $6.20\pm 0.31$
   + fine-tune 13.84 $92.66\pm 0.08$ 38.34 $43.71\pm 0.27$ 39.75 $11.04\pm 0.64$
MetaClusterKAN 13.84 $96.06\pm 0.10$ 38.34 $45.47\pm 0.44$ 39.75 $18.97\pm 0.15$
   + fine-tune 13.84 $96.41\pm 0.10$ 38.34 $45.95\pm 0.33$ 39.75 $19.16\pm 0.17$
GPTVQ 13.84 $56.29\pm 2.75$ 38.34 $24.27\pm 1.59$ 39.75 $6.07\pm 0.25$
   + fine-tune 13.84 $93.88\pm 0.25$ 38.34 $38.89\pm 0.53$ 39.75 $8.96\pm 0.18$
FastKAN 900.56 $96.78\pm 0.08$ 2,676.83 $49.10\pm 0.29$ 2,778.08 $20.56\pm 0.09$
MetaFastKAN 108.09 $95.51\pm 0.12$ 316.35 $47.40\pm 0.46$ 327.60 $19.14\pm 0.40$
ClusterFastKAN 20.54 $32.34\pm 4.16$ 57.30 $24.35\pm 0.83$ 58.71 $6.72\pm 0.41$
   + fine-tune 20.54 $86.24\pm 1.08$ 57.30 $38.89\pm 0.35$ 58.71 $11.35\pm 0.34$
MetaClusterFastKAN 20.54 $94.87\pm 0.29$ 57.30 $46.96\pm 0.44$ 58.71 $18.66\pm 0.48$
   + fine-tune 20.54 $95.59\pm 0.18$ 57.30 $47.60\pm 0.37$ 58.71 $19.05\pm 0.47$
GPTVQ 20.54 $57.62\pm 1.39$ 57.30 $20.84\pm 0.71$ 58.71 $4.34\pm 0.15$
   + fine-tune 20.54 $92.98\pm 0.35$ 57.30 $39.12\pm 0.57$ 58.71 $9.24\pm 0.37$
GramKAN 497.46 $96.76\pm 0.08$ 1,477.48 $49.57\pm 0.43$ 1,534.43 $21.89\pm 0.09$
MetaGramKAN 101.49 $95.48\pm 0.12$ 297.49 $48.24\pm 0.19$ 309.45 $19.11\pm 0.24$
ClusterGramKAN 13.96 $68.72\pm 1.65$ 38.46 $31.13\pm 0.64$ 40.58 $6.34\pm 0.43$
   + fine-tune 13.96 $95.40\pm 0.15$ 38.46 $45.81\pm 0.66$ 40.58 $13.98\pm 0.44$
MetaClusterGramKAN 14.15 $93.80\pm 0.72$ 38.65 $47.66\pm 0.25$ 40.76 $18.18\pm 0.20$
   + fine-tune 14.15 $95.32\pm 0.12$ 38.65 $48.47\pm 0.30$ 40.76 $18.96\pm 0.25$
GPTVQ 13.96 $68.35\pm 2.44$ 38.46 $16.85\pm 1.63$ 40.58 $2.36\pm 0.28$
   + fine-tune 13.96 $91.29\pm 0.47$ 38.46 $29.47\pm 1.69$ 40.58 $6.59\pm 0.61$
Table 1 demonstrates the high classification accuracy and compression factor of fully-connected MetaCluster models across all datasets and basis functions. For example, on Cifar-10, MetaClusterKAN achieves a compression factor of up to $10.7\times$ (from 410.82KB to 38.34KB) compared with MetaKAN and up to $79.9\times$ (from 3064.95 KB to 38.34KB) compared with KAN. Furthermore, it achieves this compression factor while matching the classification accuracy of MetaKAN. Although the Meta variants can achieve high accuracy without fine-tuning, this is not the case for the non-Meta variants. For example, in the case of GramKAN, it experiences a decrease in classification accuracy of up to $3.45\times$ (from 21.89% to 6.34%), which, when recovered with fine-tuning, remains a $1.57\times$ decrease in accuracy (from 21.89% to 13.98 %). Such results demonstrate the importance of leveraging a meta-learner to learn a set of activations that lie in a lower-dimensional subspace to improve classification accuracy after clustering in fully-connected models. Table 1 also compares MetaCluster models against GPTVQ (Van Baalen et al., 2024), a popular vector quantization technique that performs hessian-aware weight sharing. From our experiments, we find that hessian-aware clustering on the KAN does not yield benefits and in some cases hurts downstream performance.
We further verify the importance of MetaCluster qualitatively by plotting edge functions and the respective centroids associated with the first layer of FastKAN and MetaFastKAN clusters in Figure 2. We provide the complete set of 16 clusters for each layer in Appendix A.3. From observing the functions associated with FastKAN and MetaFastKAN, it is apparent that those of MetaFastKAN are much closer to their respective centroids than those of FastKAN. Such a result once again demonstrates the importance of using meta-learners to find a lower-dimensional functional space before clustering.
4.1.2 Convolutional Image Classification
@@
KANConv 13,634.00 $99.46\pm 0.02$ 13,654.27 $70.66\pm 2.89$ 13,744.62 $23.99\pm 0.98$
MetaKANConv 3,048.40 $99.32\pm 0.05$ 3,052.90 $68.36\pm 2.84$ 3,143.25 $35.60\pm 1.78$
ClusterKANConv 429.95 $92.16\pm 1.24$ 430.51 $33.26\pm 1.30$ 520.87 $10.81\pm 0.54$
   + fine-tune 429.95 $98.89\pm 0.09$ 430.51 $61.80\pm 0.34$ 520.87 $22.65\pm 0.81$
MetaClusterKANConv 434.77 $99.29\pm 0.06$ 435.34 $62.48\pm 2.43$ 525.70 $29.62\pm 1.65$
   + fine-tune 434.77 $99.35\pm 0.05$ 435.34 $66.85\pm 3.11$ 525.70 $35.11\pm 1.79$
FastKANConv 13,632.09 $99.11\pm 0.09$ 13,652.37 $76.30\pm 0.11$ 13,742.72 $46.36\pm 0.38$
MetaFastKANConv 3,041.57 $99.00\pm 0.11$ 3,046.08 $69.06\pm 1.15$ 3,136.43 $39.60\pm 1.13$
ClusterFastKANConv 428.03 $62.61\pm 2.70$ 428.61 $26.01\pm 3.55$ 518.97 $5.16\pm 0.31$
   + fine-tune 428.03 $96.10\pm 0.09$ 428.61 $57.45\pm 1.53$ 518.97 $20.15\pm 0.70$
MetaClusterFastKANConv 427.97 $99.01\pm 0.11$ 428.55 $68.72\pm 1.06$ 518.91 $38.75\pm 1.04$
   + fine-tune 427.97 $99.08\pm 0.06$ 428.55 $69.10\pm 1.10$ 518.91 $39.26\pm 1.10$
GramKANConv 7,586.47 $99.40\pm 0.06$ 7,597.72 $78.56\pm 1.14$ 7,688.07 $48.25\pm 0.59$
MetaGramKANConv 3,047.93 $99.44\pm 0.05$ 3,052.43 $80.91\pm 0.41$ 3,142.79 $52.07\pm 0.63$
ClusterGramKANConv 418.93 $61.17\pm 9.40$ 419.49 $12.08\pm 0.66$ 509.85 $1.77\pm 0.29$
   + fine-tune 418.93 $99.27\pm 0.08$ 419.49 $75.74\pm 0.63$ 509.85 $44.01\pm 0.42$
MetaClusterGramKANConv 418.81 $99.41\pm 0.06$ 419.38 $79.25\pm 0.59$ 509.73 $49.70\pm 0.51$
   + fine-tune 418.81 $99.47\pm 0.02$ 419.38 $80.85\pm 0.41$ 509.73 $51.20\pm 0.67$
Table 3: Impact of basis coefficient count on memory (KB) footprint and classification accuracy.
Coeff Memory Acc% Memory Acc% %Chg Memory Acc% Memory Acc% %Chg
5 316.35 $47.40\pm 0.46$ 57.30 $47.60\pm 0.37$ 0.42 3,046.08 $69.97\pm 0.62$ 428.55 $69.73\pm 0.63$ -0.34
8 316.76 $46.29\pm 0.19$ 57.68 $46.30\pm 0.24$ 0.02 3,046.52 $68.89\pm 0.63$ 440.61 $68.92\pm 0.60$ 0.05
10 317.03 $46.63\pm 0.23$ 57.93 $47.08\pm 0.27$ 0.95 3,046.81 $68.23\pm 0.30$ 448.64 $68.22\pm 0.27$ -0.01
15 317.71 $45.26\pm 0.52$ 58.55 $45.54\pm 0.44$ 0.62 3,047.54 $64.87\pm 0.63$ 468.72 $64.67\pm 0.55$ -0.30
20 318.40 $46.11\pm 0.20$ 59.18 $46.03\pm 0.27$ -0.17 3,048.26 $61.16\pm 0.72$ 488.80 $60.90\pm 0.75$ -0.43
Table 2 demonstrates that, aside from MetaFastKANConv, the Meta variants match or exceed the classification accuracy of all the non-Meta variants at a reduced memory cost. For example, in the case of MetaKANConv on Cifar-10, the reduced memory cost is up to $4.5\times$ (from 13.7 MB to 3.1 MB) compared to KANConv. Upon clustering, we find that the compression factor increases by a factor of $7.1\times$ (from 3.1 MB to 435.34 KB). In total, compared to the KANConv, the MetaClusterKANConv achieves a reduction of up to $31.7\times$ (from 13.7 MB to 435.34 KB). For such a reduction in storage cost, after fine-tuning, we find there is a negligible impact on the downstream classification accuracy when comparing KANConv to MetaKANConv. Although the non-Meta variants recover accuracy from the fine-tuning step, the recovery still leaves them far behind their original state. For instance, upon clustering, FastKANConv experiences up to a $2.93\times$ decrease (from 76.30% to 26.01 % ) in classification accuracy. Then, when recovered with fine-tuning, it still retains a $1.32\times$ decrease (from 76.30% to 57.45 %) in classification accuracy. These results again demonstrate the importance of MetaCluster for maintaining accuracy at high compression factors.
@@ 4.2.1 Basis Coefficient Count
We investigate the impact of the basis coefficient count (i.e., the number of radial basis functions) on downstream clustering performance. Since our technique employs a meta-learner to learn a lower-dimensional subspace, our approach should be resilient to changes in the basis coefficients we cluster. From Table 3, we can see that the initial accuracy of MetaFastKAN and MetaFastKANConv accuracy peaks at a coefficient count of 5 before dropping. However, even with these variations in accuracy for the coefficient count, the relative percentage change remains the same. This indicates that, regardless of the number of coefficients, our MetaCluster framework can learn a lower-dimensional manifold, facilitating easier clustering of weights.
@@ 4.2.2 Cluster Count
We verify that the MetaCluster framework scales well with the cluster count by comparing its cluster scaling properties with those of a non-meta KAN. As shown in Figure 3(b), MetaCluster maintains accuracy well as clusters are reduced, especially after fine-tuning. For the convolutional model, decreasing the cluster count from 256 to 32 lowers MetaClusterFastKANConv accuracy by only $2\%$, whereas ClusterFastKANConv drops by $21.39\%$ over the same range. Such results demonstrate our approach can achieve even greater compression, albeit with a minor loss in accuracy.
@@ 4.2.3 Meta-learner Embedding Size
The embedding size of the meta-learner directly influences the ease of clustering the meta-learner generated weights. The general trend is that as the embedding dimension increases, it becomes increasingly difficult to cluster the generated weights appropriately. While we have qualitatively demonstrated this in Figure 1, where increasing the embedding dimension makes the plot less structured, we provide a quantitative analysis of this effect in Appendix A.2.2.
4.3 Computational Cost Results
Table 4: Training/fine-tuning and inference times (in seconds)
Model MNIST CIFAR-10 CIFAR-100
Train Inference Train Inference Train Inference
KAN $2.155\pm 0.047$ $0.320\pm 0.005$ $2.661\pm 0.049$ $0.510\pm 0.008$ $2.730\pm 0.060$ $0.509\pm 0.007$
MetaKAN $2.501\pm 0.055$ $0.337\pm 0.003$ $2.963\pm 0.056$ $0.538\pm 0.012$ $2.956\pm 0.096$ $0.526\pm 0.008$
MetaClusterKAN $2.284\pm 0.048$ $0.323\pm 0.002$ $3.259\pm 0.045$ $0.521\pm 0.008$ $3.252\pm 0.091$ $0.542\pm 0.019$
KANConv $9.158\pm 0.184$ $0.922\pm 0.009$ $0.105\pm 0.085 1.1$$130\pm 0.003$ $10.370\pm 0.162$ $1.204\pm 0.023$
MetaKANConv $10.039\pm 0.228$ $0.930\pm 0.009$ $1.024\pm 0.080 1.1$$184\pm 0.025$ $10.946\pm 0.097$ $1.223\pm 0.011$
MetaClusterKANConv $9.931\pm 0.160$ $0.951\pm 0.017$ $0.605\pm 0.144 1.1$$149\pm 0.032$ $10.929\pm 0.156$ $1.202\pm 0.004$
Table 5: Clustering times (in seconds)
Model MNIST CIFAR-10 CIFAR-100
KAN $029\pm 0.011 0.0$$59\pm 0.007 0.0$$080\pm 0.022$
MetaKAN $024\pm 0.012 0.0$$48\pm 0.011 0.0$$062\pm 0.011$
KANConv $3.587\pm 0.145$ $3.650\pm 0.098$ $3.542\pm 0.114$
MetaKANConv $2.806\pm 0.072$ $2.812\pm 0.067$ $2.791\pm 0.092$
As our approach is a three-stage framework consisting of a training stage, a clustering stage, and a fine-tuning stage, each stage will contribute to the computational cost during training. Furthermore, because our approach stores the KAN architecture as a codebook and indices that index the codebook, the forward inference cost also incurs computational overhead. We quantify the clustering stage cost in Table 5 and training/fine-tuning and inference costs in Table 4. Each metric is reported in seconds, averaged over 10 runs, with standard error, on an RTX 5090 GPU. The training and inference times are reported as the time to complete one epoch and the time to evaluate on the test set, respectively.
From observing Table 5 we can see that compared to training times the clustering times are negligible (e.g., 0.029 s once vs 2.155 s per epoch). We can also see that the time required to train MetaKAN per epoch is slightly longer than KAN. Similarly, during fine-tuning (which runs for much fewer epochs than the training stage), the time per epoch is slightly longer for MetaClusterKAN (an identical architecture to ClusterKAN) than for KAN. Such a trend deviates in inference, where the MetaClusterKAN is as fast as KAN. Similar trends are seen with the convolutional models. Such results show there is minimal overhead from MetaCluster.
4.4 Equation Modeling Results
Our high-dimensional equation modeling results examined MetaCluster’s ability to model 1000-dimensional equations bounded on [1,1][-1,1]. All KAN models used were identical to our fully-connected image classification experiments. We explored the equations $f_{1}(x)=\exp\left(\frac{1}{n}\sum\sin^{2}\left(\frac{\pi x}{2}\right)\right)$, $f_{2}(x)=\sum x^{2}+x^{3}$, $f_{3}(x)=\exp\left(-\frac{1}{n}\sum x^{2}\right)$ following a similar procedure to Zhao et al. (2025). The clustered models used a cluster size of 4. Each reported MSE result is an average over 10 runs. We provide the complete set of hyperparameters in Appendix A.1.3.
Table 6: High-dimensional equation modeling MSE and memory (KB) footprint comparison.
Model Memory $f_{1}(x)$ MSE $f_{2}(x)$ MSE $f_{3}(x)$ MSE
KAN 1303.6 $5.06\times 10^{-5}$ $2.70\times 10^{-1}$ $5.30\times 10^{-7}$
MetaKAN 181.6 $3.05\times 10^{-5}$ $2.41\times 10^{0}$ $2.76\times 10^{-6}$
ClusterKAN 10.5 $9.85\times 10^{-1}$ $1.66\times 10^{2}$ $1.27\times 10^{-1}$
   + fine-tune 10.5 $2.24\times 10^{-4}$ $8.92\times 10^{-1}$ $7.82\times 10^{-6}$
MetaClusterKAN 10.5 $9.03\times 10^{-4}$ $1.58\times 10^{0}$ $3.75\times 10^{-4}$
   + fine-tune 10.5 $2.70\times 10^{-5}$ $5.04\times 10^{-1}$ $7.55\times 10^{-7}$
From Table 6, we can see that MetaClusterKAN achieves identical or, in some cases, reduced mean squared error (MSE) than the original KAN model after clustering. MetaClusterKAN can achieve these equation modeling capabilities with a 124.1x reduction in memory. This is not the case for ClusterKAN, which shows a significant increase in MSE after clustering (from $5.061\times 10^{-5}$ to $2.239\times 10^{-4}$). These results demonstrate that the principles from MetaCluster extend beyond image classification into the equation modeling domain.
@@ 5.1 Kolmogorov-Arnold Networks
KANs have evolved rapidly, with numerous variants exploring different basis functions such as B-splines, radial basis functions (RBFs), Chebyshev and Legendre polynomials, Gram polynomials, wavelets, and rational functions (Liu et al., 2024b; Li, 2024; SS et al., 2024; Bozorgasl and Chen, 2024; Aghaei, 2024). These architectures have demonstrated strong performance in scientific computing and equation modeling (Li et al., 2025; Wang et al., 2025; Coffman and Chen, 2025; Koenig et al., 2024) and have recently extended to computer vision through designs such as the Kolmogorov-Arnold Transformer (Yang and Wang, 2024; Raffel and Chen, 2025). Despite these successes, KANs face persistent challenges, including training instability, computational overhead, and a substantial increase in parameter count compared to MLPs (Yu et al., 2024; Chen et al., 2024). Our work directly addresses the memory-scaling bottleneck of KANs by creating topologically identical KANs that are smaller than comparable MLPs while maintaining accuracy.